A "bit" is a computer term for one of the digits 0 or 1. A "bit-string" of length 9 is a sequence of 9 digits, all of which are either 0 and 1.
a) How many bit strings of length 9 are there?
b) How many bit strings are between 8 and 9 bits long?

Answer by swincher4391(1107)

a) There 2*2*2*2*2*2 ... 9 times = 2^9 = 512 bit strings of length 9.
b) There are 2^8 bit strings of length 8, 2^9 of length 9. So there are 2^8 + 2^9 = 768 bit strings of length 8 or 9.